# Human-machine social systems

### v2

@@ Introduction
The emerging field of hybrid intelligence suggests that algorithms introduce novel solutions, but these may be too unfamiliar for humans to adopt. Nevertheless, machine diversity and competition might inspire alternative forms of human creativity and innovation. Developing methods to effectively combine human and machine solutions could further improve collective intelligence 165.
Box 1: Competition in high-frequency trading markets
High-frequency trading (HFT) algorithms constitute automated scripts that rely on high-speed, large-volume transactions to exploit mispricings or market signals before they disappear or are incorporated into the price 166. The phenomenon started in the mid-90s and has since spread to dominate foreign equities, foreign exchange, commodities, futures, and stock markets globally 117.
HFT algorithms process large amounts of trade history data and current news to make decisions and are thus considered the better “informed” traders 167. Some of the algorithms appear to anticipate the market, and their trades consistently predict future order flow by human traders 168. However, since most algorithms react similarly to the same public information, they exhibit less diverse trading strategies and more correlated actions among themselves compared to humans 169. Thus, although their behavior generally improves market efficiency, it can also trigger behavioral cascades and instability 170.
HFT algorithms generally act as market makers, increasing trading opportunities, reducing transaction costs, connecting buyers and sellers across venues, and submitting significant volumes of price quotes 171, 117. They facilitate price efficiency by trading in the direction of permanent price changes but opposite temporary price errors 167. This regularly acts as a stabilizing force, reducing short-term volatility 169, 172, 166. On a longer time scale, however, HFT algorithms may decease market quality by increasing volatility 173 and uncertainty 37, and by reducing trading strategy diversity 174. For instance, although the algorithms did not cause the 2010 flash crash, they exacerbated it by amplifying the volatility 175. This has led to recent efforts to regulate the speed of trading in markets, for example, by processing trades in batches at slower intervals to diminish the advantage that HFT algorithms have 176.
Box 2: Contagion on Twitter
Social bots on the micro-blogging platform Twitter (re-branded as X in 2022) are covert automated accounts designed to impersonate humans to boost followers, disseminate information, and promote products. Bots and bot detection methods have co-evolved, resulting in increasingly more sophisticated imitation or detection strategies 177, 178, 179, 180, 39, 181, 182, but detection is inherently limited due to the overlap between covert autonomous bots, managed user accounts, hacked accounts, cyborgs, sock-puppets, and coordinated botnets 24, 183, 22, 184. Estimates suggest that 9-15% of Twitter users are bots 185, 182, with bot activity typically increasing around controversial political events 186.
Twitter social bots, who do not follow social instincts but neither succumb to fatigue, engage less in social interactions via replies and mentions than humans but produce more content 187. The bots mainly retweet – a passive strategy to indicate support and gain followers – but are less successful in attracting friends and followers than humans 182. Overall, they are less connected and bot-bot (2%), bot-human (19%), and human-bot (3%) interactions are considerably less common than human-human interactions (76%) 186.
Despite their rudimentary social behavior and weak network integration, Twitter bots significantly influence political communication, public opinion, elections, and markets. They play an important role in misinformation dissemination in relation to political events 188, 189, 190, 191, 192, 193, 194, COVID-19 195, 196, and stock market investment 197. Bots can affect human interaction networks by encouraging followings and conversations 198 and amplify low-credibility content early on by targeting influential humans through replies and mentions 192. Bots’ large numbers enhance their visibility and influence to trigger deep information cascades 199. Bots equally link to true and false news from low-credibility sites, but people prefer false content, making humans ultimately responsible for the spread of false news 200.
Twitter bots significantly contribute to negative sentiment and conflict escalation. Acting from the periphery, they target central human users to exert indirect influence. They amplify existing negative sentiment and selectively promote inflammatory content, often targeting only one of the factions 186. Their success stems from exploiting human tendencies to connect with similar others and engage with messages that reinforce their beliefs 201. Consequently, bots increase ideological polarization and negatively affect democratic discourse on social media, as seen in the 2016 US presidential election 202, the 2016 UK Brexit Referendum 201, and the 2017 Catalan independence referendum 186.
In sum, Twitter’s covert social bots are considered harmful, prompting the platform to cull them 203, 204. Their strength lies in indirect action: they skew the platform’s recommendation system to bias content popularity 137 and exploit human behavioral weaknesses like attention seeking, confirmation bias, moral outrage, and ideological homophily.
Box 3: Cooperation and coordination on Wikipedia
Wikipedia, the largest and most popular free-content online encyclopedia, hosts an ecology of bots that generate articles, fix errors on pages, link to other sites and databases, tag articles in categories, identify vandals, notify users, and so on 205, 206, 207. These bots are open-source, documented, approved, registered, and tagged 208, 205. They are not sophisticated: most use basic regular expressions or straightforward heuristics, and only some incorporate machine learning techniques. They are significantly less numerous than human editors but complete a disproportionately large volume of all edits 209, 36, 207. Compared to human-human interactions, bot-bot interactions are more reciprocal and balanced but do not exhibit status effects 36. While bots are more likely to be involved in back-and-forth reverts with each other over long periods, these accidental encounters rarely indicate direct opinion conflict but constitute routine productive maintenance work or reflect conflicts existing between their human owners 210. Human editors interact mainly with policing bots, primarily by criticizing the legitimacy of the norms they enforce, rather than the sanctions themselves, suggesting that editors perceive bots as extensions of their human owners rather than independent agents 211.
Bots have been invaluable to the maintenance and operation of Wikipedia. The diversity of the bot ecology guarantees the system’s robustness and resilience. For instance, during the random outage of the anti-vandalism ClueBot NG, the website eventually caught up, albeit more slowly than usual, thanks to the heterogeneity of the quality control network, comprising instantaneous fully automated robots, rapid tool-assisted humans (cyborgs), humans editing via web browsers, and idiosyncratic batch scripts 212. Wikipedia demonstrates that successful bot governance and regulation does not have to sacrifice distributed development and diversity. The algorithmic simplicity, independence, and heterogeneity of the machines facilitate the system’s success and resilience overall but may also introduce unexpected complexities and uncertainties in communication at smaller scales 213.
Box 4: Cooperation and contagion on Reddit
Reddit is a popular news aggregation, content rating, and discussion website founded in 2005. Bots on Reddit provide internal moderation and communication, augment functionality (e.g., for mobile users), or post content, ranging from comic and playful posts by evident automated accounts such as haiku_robot and ObamaRobot, to trolling and provocative comments by undercover social bots 214. Similarly to Wikipedia, Reddit has developed norms and protocols for deploying bots 215, but similarly to Twitter, it has no effective service limitations to prevent covert and malicious automated accounts. Nevertheless, Reddit differs from Twitter in several crucial ways – content on the site is posted within communities, heavily moderated, up-/downvoted, and extensively discussed. Due to these structural differences, political misinformation, polarization, and conflict are less pronounced on the platform.
Reddit offers evidence for collaboration and contagion between humans and bots. Content moderators have widely adopted Automod – a bot that is flexible to updates, adaptable to community rules, and interpretable. The bot aids with menial tasks but does not necessarily decrease workloads as it requires continuous updates in response to changes in user behavior and language and involves high volumes of correspondence with incorrectly banned users 216. On the other side, regular users engage with evident entertainment bots, and their direct replies imitate the sentiment and the words of the bot posts 217. Thus, emotional contagion and lexical entrainment can occur between humans and bots, even when humans are aware of the simple automated script behind the bot.
Discussion
@@
The algorithms in current human-machine social systems are relatively simple. Few use sophisticated machine learning or AI, and typically, these guide narrow and specific behaviors 218, 219. Except for malicious social bots and customer-service chatbots, most machines do not mimic human qualities. Most machines are superhuman — processing vast data amounts, acting swiftly, and handling tedious tasks— or candidly non-human— resisting peer influence, not reciprocating, and acting randomly. There is a clear distinction between covert and overt bots: covert bots are more problematic than bots declaring their identity and following norms and regulations.
@@
Existing research is often biased towards engineering and optimization, lacking deeper insights from a social science perspective. The time for a new sociology of humans and machines is critical, before AI becomes more sophisticated: generative AI exhibits emergent behavior that itself requires explanation 220, 221, complicating the understanding of system dynamics.
@@
Researchers would benefit from an agent-based modeling framework that outlines distinctions between human and bot agents: utility function, optimization ability, access to information, learning, innovation/creativity, accuracy, etc. The framework could borrow concepts from other two-agent systems, such as predator–prey, principal–agent, and common pool resource models. Controlled experiments should explicitly compare human-machine, human-only and machine-only networks, and known bots against covert bots. Experiments could manipulate participants’ perceptions of algorithms’ technical specifications, agenthood 222, emotional capability, and biases. Field interventions in online communities with endemic bot populations present another promising direction. Existing examples include social bots that gain influence by engaging human users 223, 224, 225, 226, trading bots that manipulate prices in cryptocurrency markets 227, political bots that promote opposing political views to decrease polarization 228, and “drifters” to measure platform bias 94. Expanding on the cases reported here, we need observational research on additional human-machine communities and contexts such as traffic systems with human-driven and driverless vehicles, online multiplayer games comprising human players, non-player characters, and cheating code, and dating markets with AI-driven chatbots 229.
@@
Finally, research with artificial agents introduces ethical problems demanding careful elaboration and mitigation. Research protocols should minimize interventions 230, possibly deploying covert bots only where they already exist, ensuring their actions are not unusual or harmful 93. Even then, bots may still face opposition from users due to privacy concerns 223. Do people perceive certain bots as inherently deceptive? Could knowledge of the bot owner and algorithm mitigate this perception?
Implications for design
@@
Humans are resilient and successful when they exhibit high levels of efficient communication, context awareness, emotional recognition and response, and ethical and cultural sensitivities. These features should be encouraged when designing social machines to build trust, ensure legal compliance, promote social harmony, enhance user satisfaction, and achieve long-term sustainability 231, 232.
@@
Since H-H, H-M, and M-M interactions differ, machines should be specifically designed for each scenario, with separate training for each interaction. This could avoid market underperformance when bargaining algorithms trained on human-only markets adapt poorly to human-machine negotiations 64, or traffic jams when driverless vehicles trained on human driving fail to properly interact with each other 233.
@@
AI design should also adopt a hierarchy of behavioral rules and conventions guiding H-M and M-M interactions in the context of H-H interactions. Isaac Asimov’s famous Three Laws of Robotics 234 which regulate M-H interactions and self-preservation—1) not harming humans, 2) obeying humans, and 3) protecting own existence, with priority given to higher order rules—could be adapted for M-M interactions, considering specific contexts, implications, and unintended consequences 235.
@@
Further, cultural context is crucial for AI design. People’s perceptions of machines vary by age, environment, personality, and geography 236, 60. Machines reflect their developers’ culture and operate in settings with specific organizational and community norms 36. Thus, AI design for self-driving cars and assistant bots should consider local driving culture and attitudes toward domestic assistants during training.
@@
Finally, the systems under review exhibit complexity in that their emergent behavior transcends a mere aggregation of individual components, displaying what is known as “network effects.” While natural complex systems often demonstrate remarkable resilience and adaptivity 237, the human-designed complex systems discussed here, albeit to varying degrees, are not inherently adaptive. To enhance resilience and robustness, AI designers should incorporate complex adaptive system principles, like negative feedback, modularity, and hierarchical organization. For example, dense networks lacking diversity and modularity are susceptible to systemic failures 238, 239, 240 and are easy to control 241. Introducing bots that increase network diversity, introduce resistance, build resilience, or incorporate negative feedback could enhance adaptability and stability. Such configurations have improved group outcomes in forecasting 242, exploration/exploitation 243, and general knowledge tasks 151. Literature on monitoring and steering complex systems can help predict critical transitions 244, 245, 15, designing safer human-machine systems across different environments and perturbations.
Implications for policy
@@
Policymakers should also anticipate the social co-evolution of machines and humans, which will inadvertently change existing institutions. Machines can cause humans to withdraw interaction: for instance, outsourcing care to robots reduces caregivers’ empathy 246. Intelligent machines are changing the transmission and creation of human culture, altering social learning dynamics, and generating new game strategies, scientific discoveries, and art forms 247. Humans must adapt to autonomous machines just as autonomous machines must learn from and adapt to humans. Finally, ethicists should address questions such as: Should all machines be equal? Should we allow status hierarchies, possibly reflecting and exacerbating existing socio-economic inequalities?
@@ Conclusion
This survey synthesizes relatively disparate literature based on agent-based models, controlled experiments, online field interventions, and observational analyses from human-computer interaction, robotics, web science, financial economics, and computational social science under a common theoretical framework: human-machine social systems. We identify common dynamics and patterns that emerge from the interactions of humans and autonomous machines regardless of the specific context, as well as peculiarities and unique problems that concrete techno-organizational and socio-cultural environments generate. Our utmost ambition is to stimulate cumulative empirically driven and mechanism-focused sociological research in the emerging, fast-evolving field of human-AI science. At stake are new and urgent social challenges such as online misinformation, market flash crashes, cybersecurity, labor market resilience, and road safety. With increasing social connectivity and accelerating developments in AI, understanding the complex interactions between humans and autonomous machines is a challenging undertaking, but one that is crucially important for a better human future.

### v3

@@ Abstract
The emerging field of hybrid intelligence suggests that algorithms introduce novel solutions, but these may be too unfamiliar for humans to adopt. Nevertheless, machine diversity and competition might inspire alternative forms of human creativity and innovation. Developing methods to effectively combine human and machine solutions could further improve collective intelligence 151.
Implications
@@
The algorithms in current human-machine social systems are relatively simple. Few use sophisticated machine learning or AI, and typically, these guide narrow and specific behaviors 152, 153. Except for malicious social bots and customer-service chatbots, most machines do not mimic human qualities. Most machines are superhuman — processing vast data amounts, acting swiftly, and handling tedious tasks— or candidly non-human— resisting peer influence, not reciprocating, and acting randomly. There is a clear distinction between covert and overt bots: covert bots are more problematic than bots declaring their identity and following norms and regulations.
@@
Existing research is often biased towards engineering and optimization, lacking deeper insights from a social science perspective. The time for a new sociology of humans and machines is critical, before AI becomes more sophisticated: generative AI exhibits emergent behavior that itself requires explanation 154, 155, complicating the understanding of system dynamics.
@@
Researchers would benefit from an agent-based modeling framework that outlines distinctions between human and bot agents: utility function, optimization ability, access to information, learning, innovation/creativity, accuracy, etc. The framework could borrow concepts from other two-agent systems, such as predator–prey, principal–agent, and common pool resource models. Controlled experiments should explicitly compare human-machine, human-only and machine-only networks, and known bots against covert bots. Experiments could manipulate participants’ perceptions of algorithms’ technical specifications, agenthood 156, emotional capability, and biases. Field interventions in online communities with endemic bot populations present another promising direction. Existing examples include social bots that gain influence by engaging human users 157, 158, 159, 160, trading bots that manipulate prices in cryptocurrency markets 161, political bots that promote opposing political views to decrease polarization 162, and “drifters” to measure platform bias 80. Expanding on the cases reported here, we need observational research on additional human-machine communities and contexts such as traffic systems with human-driven and driverless vehicles, online multiplayer games comprising human players, non-player characters, and cheating code, and dating markets with AI-driven chatbots 163.
@@
Finally, research with artificial agents introduces ethical problems demanding careful elaboration and mitigation. Research protocols should minimize interventions 164, possibly deploying covert bots only where they already exist, ensuring their actions are not unusual or harmful 79. Even then, bots may still face opposition from users due to privacy concerns 157. Do people perceive certain bots as inherently deceptive? Could knowledge of the bot owner and algorithm mitigate this perception?
Design
@@
Humans are resilient and successful when they exhibit high levels of efficient communication, context awareness, emotional recognition and response, and ethical and cultural sensitivities. These features should be encouraged when designing social machines to build trust, ensure legal compliance, promote social harmony, enhance user satisfaction, and achieve long-term sustainability 165, 166.
@@
Since H-H, H-M, and M-M interactions differ, machines should be specifically designed for each scenario, with separate training for each interaction. This could avoid market underperformance when bargaining algorithms trained on human-only markets adapt poorly to human-machine negotiations 50, or traffic jams when driverless vehicles trained on human driving fail to properly interact with each other 167.
@@
AI design should also adopt a hierarchy of behavioral rules and conventions guiding H-M and M-M interactions in the context of H-H interactions. Isaac Asimov’s famous Three Laws of Robotics 168 which regulate M-H interactions and self-preservation—1) not harming humans, 2) obeying humans, and 3) protecting own existence, with priority given to higher order rules—could be adapted for M-M interactions, considering specific contexts, implications, and unintended consequences 169.
@@
Further, cultural context is crucial for AI design. People’s perceptions of machines vary by age, environment, personality, and geography 170, 46. Machines reflect their developers’ culture and operate in settings with specific organizational and community norms 22. Thus, AI design for self-driving cars and assistant bots should consider local driving culture and attitudes toward domestic assistants during training.
@@
Finally, the systems we reviewed exhibit complexity in that their emergent behavior transcends a mere aggregation of individual components, displaying what is known as “network effects.” While natural complex systems often demonstrate remarkable resilience and adaptivity 171, the human-designed complex systems discussed here, albeit to varying degrees, are not inherently adaptive. To enhance resilience and robustness, AI designers should incorporate complex adaptive system principles, like negative feedback, modularity, and hierarchical organization. For example, dense networks lacking diversity and modularity are susceptible to systemic failures 172, 173, 174 and are easy to control 175. Introducing bots that increase network diversity, introduce resistance, build resilience, or incorporate negative feedback could enhance adaptability and stability. Such configurations have improved group outcomes in forecasting 176, exploration/exploitation 177, and general knowledge tasks 137. Literature on monitoring and steering complex systems can help predict critical transitions 178, 179, 15, designing safer human-machine systems across different environments and perturbations.
Policy
@@
Policymakers should also anticipate the social co-development of machines and humans, which will inadvertently change existing institutions. Machines can cause humans to withdraw interaction: for instance, outsourcing care to robots reduces caregivers’ empathy 180. Intelligent machines are changing the transmission and creation of human culture, altering social learning dynamics, and generating new game strategies, scientific discoveries, and art forms 181. Humans must adapt to intelligent machines just as intelligent machines must learn from and adapt to humans. Finally, ethicists should address questions such as: Should all machines be equal? Should we allow status hierarchies, possibly reflecting and exacerbating existing socio-economic inequalities?
@@ Conclusion
This Perspective synthesizes relatively disparate literature based on agent-based models, controlled experiments, online field interventions, and observational analyses from human-computer interaction, robotics, web science, financial economics, and computational social science under a common theoretical framework: human-machine social systems. We identify common dynamics and patterns that emerge from the interactions of humans and intelligent machines regardless of the specific context, as well as peculiarities and unique problems that concrete techno-organizational and socio-cultural environments generate. Our utmost ambition is to stimulate cumulative empirically driven and mechanism-focused sociological research in the emerging, fast-evolving field of human-AI science. At stake are new and urgent social challenges such as online misinformation, market flash crashes, cybersecurity, labor market resilience, and road safety. With increasing social connectivity and accelerating developments in AI, understanding the complex interactions between humans and intelligent machines is a challenging undertaking, but one that is crucially important for a better human future.